# GTransPDM: A Graph-embedded Transformer with Positional Decoupling for Pedestrian Crossing Intention Prediction

### v1

@@ Abstract
Understanding and predicting pedestrian crossing behavioral intention is crucial for autonomous vehicles driving safety. Nonetheless, challenges emerge when using promising images or environmental context masks to extract various factors for time-series network modeling, causing pre-processing errors or a loss in efficiency. Typically, pedestrian positions captured by onboard cameras are often distorted and do not accurately reflect their actual movements. To address these issues, GTransPDM – a Graph-embedded Transformer with a Position Decoupling Module – was developed for pedestrian crossing intention prediction by leveraging multi-modal features. First, a positional decoupling module was proposed to decompose the pedestrian lateral movement and simulate depth variations in the image view. Then, a graph-embedded Transformer was designed to capture the spatial-temporal dynamics of human pose skeletons, integrating essential factors such as position, skeleton, and ego-vehicle motion. Experimental results indicate that the proposed method achieves 92% accuracy on the PIE dataset and 87% accuracy on the JAAD dataset, with a processing speed of 0.05ms. It outperforms the state-of-the-art in comparison.
@@ I INTRODUCTION
Pedestrians, as vulnerable road users, are a crucial concern in the safe driving of autonomous vehicles (AVs). Accurately understanding and predicting pedestrian behavioral intentions is essential for AVs to make proactive safety decisions [1, 2]. However, pedestrian crossing intention prediction (PCIP) continues to be a challenge in time series classification, where the input comprises a sequence of features derived from sensor data, and the output indicates the probability that a person crosses the road after a given prediction horizon [3], as shown in Fig. 1.
In existing research, RNN and Transformer based approaches are employed to develop time series models with graph neural networks (GNNs) either independently or in conjunction to represent skeletal information [4, 5, 6, 7, 8]. However, the role of skeleton information within the feature fusion paradigm remains controversial in current studies, underscoring the necessity for precise extraction of pose dynamics[9, 10]. To gain a deeper insight into pedestrian crossing behavior, various factors have been explored in existing approaches, showcasing improved performance by incorporating environmental context [11, 12, 13]. However, these methods may lead to errors in image processing, instance mask inference, or become computationally intensive.
@@
Pedestrian trajectories provide valuable insights into pedestrian behavior and have been extensively utilized in PCIP [14, 15, 16, 17]. Yet, since pedestrian positions in the image are relative to the ego-vehicle, this can hinder the models’ ability to accurately identify pedestrian movements in real-world scenarios. Typically, a pedestrian’s lateral shift toward the road is a key indicator of a potential crossing. However, the position may be distorted due to the moving ego-vehicle, making it challenging to pinpoint their actual movements. Moreover, the distance between pedestrians and the ego-vehicle greatly influences the crossing decision [18, 19]. This factor, though critical, has often been neglected in current studies due to the complexity of depth estimation.
@@
To address the above issues, a Graph-embedded Transformer with a Position Decoupling Module (GTransPDM) for PCIP was developed by merging multi-modal features. First, the PDM was introduced to represent pedestrian lateral movements using road-boundary-like reference lines. Then, the area ratio of the bounding boxes was utilized in the PDM to simulate the depth variation between pedestrian and ego-vehicle. Finally, graph convolutional blocks with learnable edge importance were combined with a Transformer-based architecture for PCIP, enabling the consideration of both spatial and temporal variations in human pose skeletons.
The main contributions are summarized as follows.
1) A fast Transformer-based multi-modal fusion framework was proposed to enhance PCIP by integrating data from various sensors.
2) A PDM was proposed to decompose the pedestrians’ lateral movements and simulate the depth variation, to eliminate the positional distortion in on-board image view.
3) A GCN-based embedding was integrated with the Transformer to capture the spatial-temporal dynamics of the human pose skeleton, allowing for accurate modeling of joint interactions and motion patterns.
II Related Work
II-A Pedestrian Intention Prediction in Time-Series Models
Understanding pedestrian crossing behavior is crucial for traffic safety. Before the rise of deep learning, machine learning (ML) methods such as Hidden Markov Models (HMM) [20], Support Vector Machines (SVM) [21, 22], and Random Forests (RF)[8] were used to identify the pedestrian’s action dynamics. However, ML relies on manual feature engineering and it is difficult to handle with high-dimensional and large-scale data, making it less effective in understanding the intricate movement patterns of pedestrians. In recent years, significant attention has been given to deep-learning-based time-series models, particularly variations of RNNs and Transformers [4, 5, 6]. However, RNNs face challenges in capturing long-range dependencies and are less efficient due to their sequential processing. To address these limitations, IntFormer first introduced Transformer encoders to model temporal dynamics [10]. PIT proposed a ViT-based [23] encoder to capture interactions between traffic agents within image frames, and stacking multiple Transformer layers to capture temporal patterns [24]. However, PIT is computationally expensive due to image processing, and IntFormer’s performance still falls short of expectations. This paper enhances these models by introducing positional decoupling and incorporating a graph-embedded Transformer for skeleton dynamics modeling, achieving both lightweight design and higher accuracy.
II-B Skeleton-Based Pedestrian Intention Prediction
Pedestrian skeletal pose offers crucial insights into actions closely associated with crossing behavior. Prior to the emergence of Graph Neural Networks (GNNs) [25], skeletal information was typically used in ML-based approaches, which depended on manually selected joint coordinates, limb angles, and head orientation [8, 14, 26]. Since pedestrian joints and limbs naturally form a graph structure, GNNs and their variants have been extensively adopted for modeling pose dynamics. Notable models include STGCN [27] and GAT [28]. In the PCIP task, skeletal information is often extracted using standalone GNNs [29, 8, 30] or combined with RNNs to capture both spatial and temporal features  [5, 9]. Motivated by the success of Transformers [31], some current studies have incorporated self-attention mechanisms into GNNs [32, 33]. However, few studies have explored the combination of GNNs and Transformers in multi-modal fusion paradigm, to the best of our knowledge.
II-C Multi-Feature Fusion for Pedestrian Intention Prediction
Pedestrian crossing behavior is affected by a variety of factors, prompting recent research to concentrate on creating feature fusion methods that incorporate elements, such as pedestrian characteristics (e.g., age, gender) [34], position  [16, 35], skeletal pose [29], and environmental factors (e.g., crosswalks, traffic lights) [36, 37]. Pedestrian position, which indicates whether a pedestrian is approaching the road, has emerged as one of the most crucial indicators [14, 15, 16, 17]. However, on-board camera views capture the relative position between the vehicle and the pedestrian, complicating the accurate identification of low-speed pedestrian movements in real-world scenarios. Moreover, the distance between the pedestrian and the ego-vehicle, a key factor in crossing behavior [18, 19], is often overlooked. Furthermore, studies that incorporate environmental context is computationally intensive [4, 5, 6, 7, 8, 24, 38]. In this study, a lightweight network with a PDM is proposed to minimize positional distortion and simulate depth variation, achieving superior performance.
III Methodology
Figure 2: Overview of the proposed framework. GTransPDM is encoded with three components: 1) A positional encoder $X_{pe}$ describes both relative position and true movement variation in on-board camera views and real world, by incorporating the proposed Positional Decoupling Module (PDM), 2) an ego-vehicle motion encoder $X_{ev}$ captures both velocity and acceleration dynamics, and 3) a GCN-based pose skeleton encoder $X_{ke}$ with learnable edge importance to modeling human action patterns. These three features are fused and passed through Transformer encoder layers for temporal modeling, followed by an MLP head for binary classification.
III-A Overview
In essence, PICP is a task for predicting action in a binary classification setting, where a sequence of observed frames serves as the input and the output is the probability of crossing after a specified time-to-event (TTE) horizon, as illustrated in Fig. 1. An overview of the proposed framework is presented in Fig. 2.
@@ III-A Overview
In the presented framework, the encoding of the network was accomplished by using position, pose, and ego-vehicle motion of the pedestrian $i$ across $T$ observation frames. Each encoder was referred to as $\mathbf{X}_{pe}$, $\mathbf{X}_{ke}$, and $\mathbf{X}_{ev}$, respectively. In positional encoder, PDM $P_{i}$ was developed to decompose true lateral movements of the pedestrian and simulate depth changes, which were then integrated with displacement $D_{i}$ and instantaneous velocity $V_{i}$ to further present the position shift in image view. Then, the ego vehicle’s motion, indicative of a potential conflict with a pedestrian, was characterized by its speed $S_{i}$ and acceleration $Acc_{i}$.
In addition, several GCN blocks equipped with residual connections were layered to capture the pedestrian pose dynamics in the spatial domain $K_{i}$. Finally, these encoded representations are combined and sent to a Transformer encoder for temporal feature extraction, followed by linking to an MLP head for classification results. As described above, the PCIP problem in this study can be outlined as follows.
@@
III-B Network Encoder
Positional Encoder. In this study, the positional encoder includes displacement $D_{i}=\left\{x_{t}-x_{0},y_{t}-y_{0}\right\}\in\mathbb{R}^{T\times 2}$, instantaneous velocity $V_{i}=\left\{x_{t}-x_{t-1},y_{t}-y_{t-1}\right\}\in\mathbb{R}^{T\times 2}$, and $P_{i}$ from the proposed PDM, where $x_{t}$ and $y_{t}$ are the absolute center coordinates of the pedestrian bounding box, and $t\in(0,...,T)$. $D_{i}$ and $V_{i}$ collectively describe the overall and frame-by-frame displacement of the pedestrian relative to the ego-vehicle.
Some previous studies have treated pedestrian position as absolute coordinates in the image view from the on-board camera [14, 15, 16]. However, when the vehicle is in motion, this can lead to positional distortion, making it difficult for the network to accurately detect pedestrians approaching the road. Moreover, depth is a key factor in crossing behavior but is rarely considered in the existing research. To this end, the PDM was proposed to address the above issue, as shown in Fig. 2 (a). Firstly, two reference lines were set to represent the road boundary coordinates $(x_{l},y_{l})$. Then, the disparity between the pedestrian center and the reference line at frame $t$ was calculated, namely $(\Delta x_{t},\Delta y_{t})$, where $\Delta x_{t}=x_{t}-x_{l}$, $\Delta y_{t}=y_{t}-y_{l}$ to indicate position relative to the line of sight. Finally, the area ratio $R$ was employed to combine with the position variations in PDM to mimic the depth shift in image view. The whole features in $P_{i}$ can be represented as:
@@ III-B Network Encoder
where: $A_{t}$ is the bounding box erea at frame $t$. $\alpha$ is a factor to magnify the value, was set as $\alpha=100$.
Each feature in $\mathbf{X}_{pe}=[P_{i},D_{i},V_{i}]$ was fed into a fully connected (FC) layer for feature extraction, then concatenated along the channel dimension for feature fusion. Finally, another FC layer was applied for linear projection to produce the output, denoted as:
@@
Ego-vehicle Motion Encoder. The motion of the ego-vehicle represents vehicles that are about to conflict with a pedestrian. Thus, speed $S_{i}=\left\{s_{t}\right\}\in\mathbb{R}^{T\times 1}$, and acceleration $Acc_{i}=\left\{acc_{t}\right\}\in\mathbb{R}^{T\times 1}$ were utilized simultaneously to describe it. To represent the global $Acc_{i}$, it was calculated as:
@@
Skeleton Pose Encoder. The pedestrian skeleton pose reflects the action highly related to the crossing behavior, which can be represented as a graph $G=\{V,E\}$ naturally, where the joints are represented as nodes $V$ and limbs as edges $E$. $N=20$ keypoints $K_{i}=\left\{K_{i}^{1},K_{i}^{2},\ldots,K_{i}^{N}\right\}\in\mathbb{R}^{T\times N\times 3}$ was utilized, each point is represented as $K_{i}^{j}=\left\{\left(x_{t}^{\prime},y_{t}^{\prime},s_{t}\right)\mid j\in 1,\ldots,N\right\}$. The adjacency matrix $\mathbf{A}$ includes the skeleton naturally connections and self-connected matrix. To avoid redundancy with position, each absolute coordinates of keypoint $(x_{t}^{k},y_{t}^{k})$ in the image was normalized with its corresponding top-left bounding box coordinates $(x_{t}^{btl},y_{t}^{btl})$ as: $\left(x_{t}^{\prime},y_{t}^{\prime}\right)=\left(x_{t}^{k}-x_{t}^{btl},y_{t}^{k}-y_{t}^{btl}\right)$, and $s_{t}$ is the confidence score.
Inspired by [39], four stacked GCN blocks with residual connections were introduced as feature extractors of $K_{i}$, as shown in Fig. 2 (b). Furthermore, learnable edge importance was also applied in the network training. The entire process can be denoted as:
$\begin{gathered}\widehat{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{AD}{}^{-\frac{1}{2}}\\ \mathbf{X}_{i}=\text{ BatchNorm }\left(K_{i}\right)\\ \mathbf{H}_{0}=\sigma\left((\mathbf{E}_{0}\odot\widehat{\mathbf{A}})\mathbf{X}_{i}\mathbf{W}_{0}\right)\\ \mathbf{H}_{1}=\sigma\left((\mathbf{E}_{1}\odot\widehat{\mathbf{A}})\mathbf{H}_{0}\mathbf{W}_{1}+\operatorname{Conv2D}\left(\mathbf{H}_{0};\mathbf{W}_{c1}\right)\right)\\ \mathbf{H}_{2}=\sigma\left((\mathbf{E}_{2}\odot\widehat{\mathbf{A}})\mathbf{H}_{1}\mathbf{W}_{2}+\operatorname{Conv2D}\left(\mathbf{H}_{1};\mathbf{W}_{c2}\right)\right)\end{gathered}$ (6)
$\mathbf{H}_{3}=\sigma\left((\mathbf{E}_{3}\odot\widehat{\mathbf{A}})\mathbf{H}_{2}\mathbf{W}_{3}+\operatorname{Conv2D}\left(\mathbf{H}_{2};\mathbf{W}_{c3}\right)\right)$
where: $\mathbf{D}$ is the diagonal node degree matrix. $\mathbf{E}_{0}\sim\mathbf{E}_{3}$ and $\mathbf{H}_{0}\sim\mathbf{H}_{3}$ represent the learnable edge importance and feature map in each layer respectively, $\sigma(\cdot)$ indicates the ReLU activate function, and $\odot$ is Hadamard product operation. Further, the last layer $\mathbf{H}3$ was flattened to $\mathbb{R}^{T\times(N\times d_{hid})}$ and linear projected as the output of $\mathbf{X}_{ke}$:
@@
where: $d_{hid}$ is the dimension of hidden layers in GCN.
III-C Feature Fusion, Transformer Encoder and Prediction
After encoder layers, the $\mathbf{X}_{pe}$,$\mathbf{X}_{ev}$,and $\mathbf{X}_{ke}$ were generated and concatenated along the channel dimension for feature fusion, with another FC layer for linear projection.
@@ III-C Feature Fusion, Transformer Encoder and Prediction
So far, $\mathbf{X}$ includes the features from all modalities. Benefiting from the multi-head attention (MHA) mechanism in long-range dependencies, Transformer encoder layers [31] were utilized to model on temporal dimension. Firstly, positional encoding (PE) was added in temporal steps:
$\begin{gathered}\mathbf{PE}(t,2i)=\sin{\left(t/10000^{2i/d_{m}}\right)}\\ \mathbf{PE}(t,2i+1)=\cos{\left(t/10000^{2i/d_{m}}\right)}\end{gathered}$ (9)
where: $i$ is the index of the embedding dimension $d_{m}$. Further, the MHA layer can be denoted as:
$\begin{gathered}\mathbf{X}^{\prime}=\mathbf{X}+\mathbf{PE}\\ \mathbf{Q}=\mathbf{X}^{\prime}\mathbf{W}_{\mathbf{Q}},\mathbf{K}=\mathbf{X}^{\prime}\mathbf{K}_{\mathbf{Q}},\mathbf{V}=\mathbf{X}^{\prime}\mathbf{V}_{\mathbf{Q}}\\ \operatorname{head}_{\boldsymbol{i}}=\operatorname{Att}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{softmax}\left(\frac{\mathbf{QK}^{\mathbf{T}}}{\sqrt{\mathbf{d}_{\mathbf{k}}}}\right)\mathbf{V}\\ \mathbf{H}_{\mathbf{hid}}=\operatorname{Concat}\left(\text{ head }_{1},\ldots,\text{ head }_{h}\right)\mathbf{W}_{\boldsymbol{O}}\\ \mathbf{X}^{\prime\prime}=\text{ LayerNorm }\left(\mathbf{X}^{\prime}+\mathbf{H}_{\text{hid }}\right)\end{gathered}$ (10)
With the feature $\mathbf{X}^{\prime\prime}$ from MHA and layer normalization, a feed-forward network (FFN) with two FC layers was employed, denoted as:
$\boldsymbol{FFN}\left(\mathbf{X}^{\prime\prime}\right)=\mathbf{FC}\left(\mathbf{FC}\left(\mathbf{X}^{\prime\prime},\mathbf{W}_{\boldsymbol{f}\mathbf{1}}\right),\mathbf{W}_{\boldsymbol{f}\mathbf{2}}\right)$ (11)
$\mathbf{X}_{\text{out }}=\operatorname{LayerNorm}\left(\mathbf{X}^{\prime\prime}+\boldsymbol{FFN}\left(\mathbf{X}^{\prime\prime}\right)\right)$
Following the block containing MHA and FFN layers above, a few blocks were sequentially connected for the entire Transformer encoder in temporal modeling. Finally, the output $\mathbf{X}_{\text{out }}$ was flattened and an MLP head with a single FC layer was employed for binary classification:
$\hat{y}=\text{Softmax}\left(\mathbf{FC}\left(\text{ Flatten }\left(\mathbf{X}_{\text{out }}\right)\right)\right)$ (12)
IV Experiments
IV-A Datasets
The effectiveness of the proposed method was evaluated using two large-scale datasets in the context of autonomous driving – JAAD and PIE. JAAD [40] includes 346 short video clips between 5-10 seconds long derived from over 240 hours of recordings collected in the urban of North America and Eastern Europe. PIE [41] contains more than 6 hours of footage captured in the downtown of Toronto, Canada. Both datasets are annotated with the crossing point in time, enabling the TTE calculation. The video frame rate from the two datasets is 30 FPS. Instead of ego-vehicle speed, five ego-vehicle motion states were used in the $\mathbf{X}_{ev}$ of JAAD, as velocity data are not available.

### v2

@@ Abstract
Understanding and predicting pedestrian crossing behavioral intention is crucial for the driving safety of autonomous vehicles. Nonetheless, challenges emerge when using promising images or environmental context masks to extract various factors for time-series network modeling, causing pre-processing errors or a loss of efficiency. Typically, pedestrian positions captured by onboard cameras are often distorted and do not accurately reflect their actual movements. To address these issues, GTransPDM – a Graph-embedded Transformer with a Position Decoupling Module – was developed for pedestrian crossing intention prediction by leveraging multi-modal features. First, a positional decoupling module was proposed to decompose pedestrian lateral motion and encode depth cues in the image view. Then, a graph-embedded Transformer was designed to capture the spatio-temporal dynamics of human pose skeletons, integrating essential factors such as position, skeleton, and ego-vehicle motion. Experimental results indicate that the proposed method achieves 92% accuracy on the PIE dataset and 87% accuracy on the JAAD dataset, with a processing speed of 0.05ms. It outperforms the state-of-the-art in comparison.
Index Terms:
Pedestrian behavior, Intention prediction, Time-series model, Traffic safety.
@@ I Introduction
Pedestrians, as vulnerable road users, are a crucial concern for the safe operation of autonomous vehicles (AVs). Accurately understanding and predicting pedestrian behavioral intentions is essential for AVs to make proactive safety decisions [1, 2]. Pedestrian crossing intention prediction (PCIP) continues to be a challenge in time series classification, where the input comprises a sequence of features derived from sensor data, and the output indicates the probability that a person will cross the road after a given prediction horizon [3] (Fig. 1).
@@
In existing research, RNN-based approaches have been employed to develop time series models, either independently or in combination with graph neural networks (GNNs), to represent skeletal information [4, 5, 6, 7, 8]. However, the role of skeleton information within the feature fusion paradigm remains controversial in current studies, underscoring the necessity for precise extraction of pose dynamics [9, 10]. To gain a deeper insight into pedestrian crossing behavior, various factors have been explored in existing approaches, showcasing improved performance by incorporating environmental context [11, 12, 13, 14]. However, these methods may lead to errors in image processing or instance mask inference, and may also become computationally intensive. Pedestrian trajectories provide valuable insights into pedestrian behavior and have been extensively utilized in PCIP [15, 16, 17, 18]. Yet, since pedestrian positions in the image are relative to the ego-vehicle, this can hinder a model’s ability to accurately identify pedestrian movements in real-world scenarios. Typically, a pedestrian’s lateral shift toward the ego-vehicle is a key indicator of a potential crossing. However, the position may be distorted due to the motion of the ego-vehicle, making it challenging to pinpoint their actual movements. Moreover, the distance between pedestrians and the ego-vehicle greatly influences the crossing decision [19, 20]. This factor, though critical, has often been neglected in current studies due to the complexity of depth estimation.
@@
To address the above issues, a Graph-embedded Transformer with a Position Decoupling Module (GTransPDM) was developed by merging multi-modal features. Based on the observation that crossing pedestrians in conflict with the ego-vehicle inevitably exhibit lateral movement toward it, we propose a simple yet intuitive PDM to enhance pedestrian lateral movement recognition using reference lines. Then, the area ratio of the bounding boxes was proposed as a proxy representation of depth variation, demonstrating strong consistency with actual depth measurements. Finally, residual graph convolutional blocks with learnable edge importance were designed within a Transformer-based architecture for PCIP, enabling the consideration of spatio-temporal variations in human pose skeletons. The main contributions are summarized as follows.
  • We are the first to address the motion disparity between ego-vehicles and pedestrians to improve pedestrian position representations from on-board sensors. Building on this insight, we introduce PDM, a method that decouples the movements of AVs and pedestrians to enhance the recognition of pedestrians approaching the vehicle. Moreover, we leverage the area ratio of bounding box as depth cues instead of costly depth estimation.
  • We propose a fast Transformer-based multi-modal fusion framework to enhance PCIP by integrating data from various sensors. Notably, we develop residual graph convolutional blocks with learnable edge importance to capture the spatio-temporal dynamics of the human pose skeleton, which allows for accurate modeling of joint interactions and motion patterns.
  • Extensive experiments on the large-scale PIE and JAAD datasets demonstrate the superior performance of our proposed method, achieving an inference time of 0.05 ms using a single GPU, outperforming existing methods.
II Methodology
Figure 2: Overview of the proposed framework.
II-A Overview
@@
An overview of the proposed framework is presented in Fig. 2. In the presented framework, the encoding of the network was performed using position, pose, and ego-vehicle motion toward pedestrian $i$ across $T$ observation frames. Each encoder was referred to as $\mathbf{X}_{pe}$, $\mathbf{X}_{ke}$, and $\mathbf{X}_{ev}$, respectively. In positional encoder, PDM $P_{i}$ was developed to decompose true lateral movements of the pedestrian and reflect depth changes, which were then integrated with displacement $D_{i}$ and instantaneous velocity $V_{i}$ to further represent the positional shift in image view. Then, the ego-vehicle’s motion, which indicates potential conflict with a pedestrian, was characterized by its speed $S_{i}$ and acceleration $Acc_{i}$. In addition, several Graph Convolutional Network (GCN) blocks equipped with residual connections were layered to capture spatial dynamics of the pedestrian pose $K_{i}$. Finally, these encoded representations were combined and fed into a Transformer encoder for temporal modeling, followed by an MLP head for classification. As described above, the PCIP problem in this study can be outlined as follows.
@@
II-B Network Architecture
Positional Encoder. In this study, the positional encoder includes displacement $D_{i}=\left\{x_{t}-x_{0},y_{t}-y_{0}\right\}\in\mathbb{R}^{T\times 2}$, instantaneous velocity $V_{i}=\left\{x_{t}-x_{t-1},y_{t}-y_{t-1}\right\}\in\mathbb{R}^{T\times 2}$, and $P_{i}$ from the proposed PDM, where $x_{t}$ and $y_{t}$ are the center coordinates of pedestrian bounding boxes, and $t\in(0,...,T)$. $D_{i}$ and $V_{i}$ collectively describe the overall and frame-by-frame displacement of the pedestrian relative to the ego-vehicle. Our proposed PDM is shown in Fig. 2 (a). Firstly, the region of interest (ROI) for the crossing behavior was enclosed using two reference lines $(x_{l},y_{l})$. Then, the disparity between the pedestrian center and the reference line at frame $t$ was calculated, namely $(\Delta x_{t},\Delta y_{t})$, where $\Delta x_{t}=x_{t}-x_{l}$, $\Delta y_{t}=y_{t}-y_{l}$ to indicate position disparity between them. Finally, the area ratio $R$ was employed to combine with the position variations in PDM to capture the depth shift in image view. The whole features in $P_{i}$ can be represented as:
@@ II-B Network Architecture
where: $A_{t}$ is the bounding box area at frame $t$. $\alpha$ is a scaling factor and was set as $\alpha=100$. Each feature in $\mathbf{X}_{pe}=[P_{i},D_{i},V_{i}]$ was passed through a fully connected (FC) layer, then concatenated along the channel dimension for fusion. Finally, another FC layer was applied for linear projection to generate the final representation:
@@
Ego-vehicle Motion Encoder. The motion of the ego-vehicle indicates a potential conflict with a pedestrian. Thus, speed $S_{i}=\left\{s_{t}\right\}\in\mathbb{R}^{T\times 1}$ and acceleration $Acc_{i}=\left\{acc_{t}\right\}\in\mathbb{R}^{T\times 1}$ of the ego-vehicle were utilized simultaneously to describe it. To represent the global $Acc_{i}$, it was calculated as:
@@
Skeleton Pose Encoder. The pedestrian skeleton pose reflects the action highly related to the crossing behavior, which can be represented as a graph $G=\{V,E\}$ naturally, where the joints are represented as nodes $V$ and limbs as edges $E$. $N=20$ keypoints $K_{i}=\left\{K_{i}^{1},K_{i}^{2},\ldots,K_{i}^{N}\right\}\in\mathbb{R}^{T\times N\times 3}$ was utilized, with each point is represented as $K_{i}^{j}=\left\{\left(x_{t}^{\prime},y_{t}^{\prime},s_{t}\right)\mid j\in 1,\ldots,N\right\}$. The adjacency matrix $\mathbf{A}$ includes the skeleton naturally connections and self-connected matrix. To avoid redundancy with position, each absolute coordinates of keypoint $(x_{t}^{k},y_{t}^{k})$ in the image was normalized with its corresponding top-left bounding box coordinates $(x_{t}^{btl},y_{t}^{btl})$ as: $\left(x_{t}^{\prime},y_{t}^{\prime}\right)=\left(x_{t}^{k}-x_{t}^{btl},y_{t}^{k}-y_{t}^{btl}\right)$, and $s_{t}$ is the confidence score. Inspired by the convolution in graph structure [21], four stacked GCN blocks with residual connections were proposed as feature extractors of $K_{i}$, as shown in Fig. 2 (b). Furthermore, learnable edge importance was also applied in the network training. The entire process can be denoted as:
$\begin{gathered}\widehat{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{AD}{}^{-\frac{1}{2}}\\ \mathbf{X}_{i}=\text{ BatchNorm }\left(K_{i}\right)\\ \mathbf{H}_{0}=\sigma\left((\mathbf{E}_{0}\odot\widehat{\mathbf{A}})\mathbf{X}_{i}\mathbf{W}_{0}\right)\\ \mathbf{H}_{l}=\sigma\left((\mathbf{E}_{l}\odot\widehat{\mathbf{A}})\mathbf{H}_{l-1}\mathbf{W}_{l}+\operatorname{Conv2D}\left(\mathbf{H}_{l-1};\mathbf{W}_{cl}\right)\right)\\ \end{gathered}$ (6)
where: $\mathbf{D}$ is the diagonal node degree matrix. $\mathbf{E}_{l}$ represents the learnable edge matrix that align with the dimension of $\mathbf{A}$ in layer $l$ ($l=1,2,3$), while $\mathbf{H}_{l}$ means the feature map. $\sigma(\cdot)$ indicates the ReLU activate function, and $\odot$ is Hadamard product operation. Further, the last layer $\mathbf{H}_{3}$ was flattened to $\mathbb{R}^{T\times(N\times d_{hid})}$ and linear projected as the output of $\mathbf{X}_{ke}$:
@@
where: $d_{hid}$ is the dimension of hidden layers in GCN. After encoder layers, the $\mathbf{X}_{pe}$,$\mathbf{X}_{ev}$,and $\mathbf{X}_{ke}$ were concatenated along the channel dimension for feature fusion, with another FC layer mapping to the Transformer embedding space.
@@
So far, $\mathbf{X}$ comprises features from all modalities. To capture long-range dependencies, Transformer encoder layers [22] were employed by leveraging the self-attention mechanism to model the temporal dimension. Additionally, positional encoding (PE) was applied to preserve temporal order. After stacking several Transformer encoder layers, the output $\mathbf{X}_{\text{out}}$ was flattened and passed through an MLP head with a single fully connected layer for binary classification.
$\hat{Y}_{i}=\text{softmax}\left(\mathbf{FC}\left(\text{ Flatten }\left(\mathbf{X}_{\text{out }}\right)\right)\right)$ (9)
III Experiments
III-A Datasets and Implementation Details